OpenStudy (anonymous):

y varies directly as the square of x. When x = 4, y = 32. Find x when y = 18.

OpenStudy (anonymous):

y = x^2 + C, we need to find C. 32 = 4^2 + C -> 32 = 16 + C -> C = 16. Plug in 18 as y, 18 = x^2 + 16, 2 = x^2. Thus x is the square root of 2. I think this is the solution unless you mean something else.

OpenStudy (anonymous):

I disagree. y varies directly with the square of x implies: y = kx^2 You need to find k. There is no other term.